A. A researcher is conducting an investigation on the income of the alumni of a certain university 5 years after graduation. The monthly income of 200 responds were taken and are presented in a distribution as shown below

Classes f

3, 500- 4, 999 6

5, 000- 6, 499 23

6, 500- 7, 999 36

8, 000- 9, 499 40

9, 500- 10,999 59

11,000- 12, 499 20

12, 500- 13,999 8

14, 000- 15, 499 6

15, 500- 16, 999 2


a. Range

b. Compute the value of the variance

c. Compute the value of standard deviation

B. The efficiency rating of 12 employees of a certain department were taken and are below.

81,86,68,69,78,93,81,83,71,88,95,83


Using the formula of ungrouped data compute the value of the following.

a. Range

b. Variance

C. Standard deviation



1
Expert's answer
2021-12-28T09:34:16-0500

A.

a. In the case of grouped data, the range is the difference between the upper boundary of the highest class and the lower boundary of the lowest class.


"Range_1=16999.5-3499.5=13500"

Range is also calculated by using the difference between the mid points of the highest class and the lowest class


"Range_2=16249.5-4249.5=12000"

b.

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 Classes & M & f & M\cdot f & M^2\cdot f \\ \hline
 3500- 4999 & 4249.5 & 6 & 25497 &108349501.5 \\
 \hdashline
 5000- 6499 & 5749.5 & 23 & 132238.5 &760305255.75 \\
 \hdashline
 6500- 7999 & 7249.5 & 36 & 260982 &1891989009 \\
 \hdashline
 8000- 9499 & 8749.5 & 40 & 349980 &3062150010 \\
 \hdashline
 9500- 10999 & 10249.5 & 59 & 604720.5 &6198082764.75 \\
 \hdashline
 11000- 12499 & 11749.5 & 20 & 234990 &2761015005\\
 \hdashline
 12500- 13999 & 13249.5 & 8 & 105996 &1404394002 \\
 \hdashline
 14000-15499 & 14749.5 & 6 & 88497 &1305286501.5 \\
 \hdashline
 15500- 16999 & 16249.5 & 2 & 32499 & 528092500.5 \\
\end{array}"

"n=200"


"mean=\dfrac{1}{n}\displaystyle\sum_{i=1}^nM_if_i"

"=\dfrac{1}{200}(25497+132238.5+260982"

"+349980+604720.5+234990"

"+105996+88497+32499)=9177"

The sample variable is calculated as follows:


"Var(X)=s^2=\dfrac{1}{n-1}(\displaystyle\sum_{i=1}^nM_i^2f_i-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nM_if_i)^2)"

"\displaystyle\sum_{i=1}^nM_i^2f_i=108349501.5+760305255.75"

"+1891989009+3062150010+6198082764.75"

"+2761015005+1404394002+1305286501.5"

"+528092500.5=18019664550"

"Var(X)=s^2"

"=\dfrac{1}{200-1}(18019664550-\dfrac{1}{200}(1835400)^2)"

"=5910546.482412"


c. The sample standard deviation is


"s=\sqrt{s^2}=2431.161550"

B.

a.


"68,69, 71,78,81,81,83,83,86,88,93,95"

"n=12"

"Range=95-68=27"

b.

"mean=\bar{x}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nx_i=\dfrac{1}{12}(68+69+71+78"

"+81+81+83+83+86+88+93+95)"

"=\dfrac{976}{12}=\dfrac{244}{3}\approx83.3333"

"Var(X)=s^2=\dfrac{1}{n-1}\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2"

"=\dfrac{1}{12-1}((68-\dfrac{244}{3})^2+(69-\dfrac{244}{3})^2"

"+(71-\dfrac{244}{3})^2+(78-\dfrac{244}{3})^2+(81-\dfrac{244}{3})^2"

"+(81-\dfrac{244}{3})^2+(83-\dfrac{244}{3})^2+(83-\dfrac{244}{3})^2"

"+(86-\dfrac{244}{3})^2+(88-\dfrac{244}{3})^2+(93-\dfrac{244}{3})^2"

"+(95-\dfrac{244}{3})^2)=\dfrac{2528}{33}\approx76.606061"

c.


"s=\sqrt{s^2}=\sqrt{\dfrac{2528}{33}}\approx8.7525"